OpenStudy (anonymous):

solve by factoring x^3 - 64=0

OpenStudy (anonymous):

x^3 - 64=0 x^3 = 64 \[x \sqrt[3]{64} = 4\]

OpenStudy (anonymous):

\[x = \sqrt[3]{64} = 4\]

OpenStudy (anonymous):

has to be by factoring you have to use SOAP then use quadratic formula

OpenStudy (anonymous):

difference of cubes... has 3 roots, 1 real and 2 complex

OpenStudy (mathmath333):

\(\tt \color{black}{\text{use formula}}\) \(\large\tt \color{black}{a^3-b^3=(a-b)(a^2+ab+b^2)}\)

OpenStudy (anonymous):

i used the formula. now i have to find the Zeros (answers)

OpenStudy (mathmath333):

\(\tt \color{black}{\text{u will get one solution x=4 thats clear}}\) \(\large \tt \color{blue}{for ~~x^2+4x+16=0}\) \(\tt \color{black}{\text{use formula}}\) \(\large \tt \color{blue}{x=\dfrac{-b\pm \sqrt{b^2-4ac}}{2a}}\)